OpenStudy (blackbird02):

Proving trigonometric identities. How to solve this using only the right side of the equation.

OpenStudy (blackbird02):

\[\sqrt{(5\cos(\theta)+12\sin(\theta))^2+(12\cos(\theta)-5\sin(\theta))^2}=13\]

OpenStudy (blackbird02):

I have already solved this by manipulating the left side but our teacher said that we could also solve it by manipulating only the right side.

OpenStudy (anonymous):

work backwards from your answer on the left side and that's your answer for the right side

OpenStudy (blackbird02):

i mean, how do you make "13" equal to the left side of the equation?

OpenStudy (anonymous):

how did you make the left side of the equation equal to 13?

OpenStudy (blackbird02):

@Auxuris first I simplified the terms \[\sqrt{(25\cos^2(\theta)+120\cos(\theta)\sin(\theta)+144\sin^2(\theta))+(144\cos^2(\theta)-120\cos(\theta)\sin(\theta)+25\sin^2(\theta)}\]

OpenStudy (blackbird02):

which became \[\sqrt{169\cos^2(\theta)+169\sin^2(\theta)}=13\] then factoring out 169, \[\sqrt{169(\cos^2(\theta)+\sin^2(\theta))}=13\] since \[\cos^2(\theta)+\sin^2(\theta)=1\] the equation will be \[\sqrt{169}=13\] and the square root of 169 is 13, so 13=13

OpenStudy (jhannybean):

That is correct.

OpenStudy (michele_laino):

@blackbird02 please, try this: \[13=13*1=13*[(\cos \theta)^{2}+(\sin \theta)^{2}]=13 (\cos \theta)^{2}+13(\sin \theta)^{2}\]

OpenStudy (blackbird02):

@Michele_Laino What do I do next?

OpenStudy (michele_laino):

Sorry, better may be this: \[13=(\sqrt{13*1})^{2}=(\sqrt{13[(\cos \theta)^{2}+(\sin \theta)^{2}]})^{2}\]

OpenStudy (michele_laino):

next substitute, like above, and square both sides of your equality

OpenStudy (michele_laino):

sorry again, \[13=\sqrt{169*1}=\sqrt{169*[(\cos \theta)^{2}+(\sin \theta)^{2}]}=...\]

OpenStudy (michele_laino):

now substitute the above quantity at the right side, and then square both sides of the resultant equality

OpenStudy (michele_laino):

@blackbird02 try please

OpenStudy (blackbird02):

Do I have to square the left side of the equation too?

OpenStudy (michele_laino):

yes! both sides, please!

OpenStudy (blackbird02):

Just a sec

OpenStudy (michele_laino):

of course, first substitute the right side, how I have indicated above please!

OpenStudy (blackbird02):

Will there be any other way with out manipulating the left side?

OpenStudy (michele_laino):

I try to find it

OpenStudy (blackbird02):

I'm just curious to know how to make the right side equal to the left side without touching or manipulating the left part of the equation. Our calculus teacher said there's a way but I don't know how.

OpenStudy (michele_laino):

maybe I found it: please try to substitute to right side of your equality this: \[13=\sqrt{169*[(\cos (\theta+\alpha))^{2}+(\sin(\theta+ \alpha))^{2}}\] where alpha is such that: \[\sin \alpha =\frac{ 5 }{ 13 }, \cos \alpha=\frac{ 12 }{ 13 }\] then square both sides of your equality without manipulating the left side

OpenStudy (michele_laino):

@blackbird02 try please, I think it is the method of your teacher